OpenStudy (anonymous):

. 6 + (-3) + 4 = 4 + 6 what property is use here

OpenStudy (anonymous):

It's hard to tell, it looks like a no solution because 3 + 4 = 7 and 6 + 4 = 10. It sort of confuses me.. :l

OpenStudy (anonymous):

i know me too

OpenStudy (anonymous):

What are the answer choices if it's a multiple choice problem?

OpenStudy (anonymous):

just wants to know what property

OpenStudy (anonymous):

and for this one too . –(7 + 8) = -7 + (-8)

OpenStudy (anonymous):

Try putting "No solution" then.

OpenStudy (anonymous):

Er. That one is associative of addition

OpenStudy (anonymous):

the answer is 42^flyingpigs obviously

OpenStudy (anonymous):

ok thxs

OpenStudy (anonymous):

-15 = -15. So –(7 + 8) = -7 + (-8) is Associative of Addition Property. And you're welcome. (;

OpenStudy (anonymous):

what is this one . (4 + 5) + 9 = 4 + (5 + 9)

OpenStudy (anonymous):

Associative of Addition also. c:

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

They equal the same answer except the numbers and parenthesis are switched around.

OpenStudy (anonymous):

And yw again. :D

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

lily is wrong im right

OpenStudy (anonymous):

Well. She's asking for the Property. There are different kinds of Properties. Commutative Property of Addition, Commutative Property of Multiplication, Associative Property of Addition, Associative Property of Multiplication, Identity Property of Addition, Identity Property of Multiplication, Property of Zero, and Distributive Property.

OpenStudy (anonymous):

i'm still right

OpenStudy (anonymous):

If you're trying to troll around with me - it isn't working. (; Now. Where's the block button..